Start time | End time | Error code | Comments | Channels |
---|---|---|---|---|
17-06-2022 08:00-00 | 17-06-2022 08:16-00 | 2 | Field Mill cleaning maintenance | PG, |
19-05-2022 10:15-00 | 19-05-2022 10:19-00 | 2 | Suntracker routine cleaning maintenance | Sd, Sb, |
15-03-2022 14:43-00 | 15-03-2022 14:48-00 | 2 | Suntracker routIne cleaning maintenance | Sd, Sb, |
25-02-2022 14:11-00 | 25-02-2022 14:14-00 | 2 | Suntracker routIne cleaning maintenance | Sd, Sb, |
28-01-2022 11:25-00 | 28-01-2022 11:32-00 | 2 | Suntracker routIne cleaning maintenance | Sd, Sb, |
29-09-2021 13:57-00 | 29-09-2021 14:14-00 | 2 | Suntracker routine cleaning maintenance | Sd, Sb, |
28-06-2021 13:30-00 | 28-06-2021 14:30-00 | 2 | Rain gauge testing | Rain, |
01-06-2021 09:52-00 | 01-06-2021 10:01-00 | 2 | Suntracker routine cleaning maintenance | Sd, Sb, |
24-05-2021 12:11-00 | 24-05-2021 12:20-00 | 1 | Tipping bucket rain gauge relocated to orginal site | Rain, |
19-05-2021 10:41-00 | 19-05-2021 11:13-00 | 2 | Suntracker routine cleaning maintenance | Sd, Sb, |
05-11-2020 09:20-00 | 11-01-2021 12:00-00 | 2 | Sonic and licor disconnected | |
18-02-2020 11:35-00 | 18-02-2020 12:35-00 | 1 | Connection boxes dismantled for repair | U2, U5, U10, U0.56, U0.8, U1.12, U1.6, U2.24, U3.2, U4.48, U6.4, Rain, |
10-02-2020 09:00-00 | 10-02-2020 10:15-00 | 2 | Disconnection and replacement of sensor | PG, |
01-07-2019 10:00-00 | 09-08-2019 09:10-00 | 2 | Disconnection of sensor for checking | PG, |
23-04-2019 13:55-00 | 23-04-2019 14:05-00 | 2 | Soil moisture test probe removed for grass cutting | SoilMoisture, |
23-04-2019 09:20-00 | 23-04-2019 10:05-00 | 2 | Sensors disconnected for installation of new connectors | Sg, Sd, Sb, |
26-03-2019 11:14-00 | 26-03-2019 12:50-00 | 2 | Sensors disconnected for re wiring during installation of Biral SWS250 sensor on post | Sg, Sd, Sb, |
13-02-2019 11:50-00 | 13-02-2019 12:00-00 | 2 | Suntracker routine cleaning maintenance | Sd, Sb, |
10-01-2018 00:00-00 | 14-01-2018 10:20-00 | 2 | Disconnection of sensor for investigation of failure | RH, |
21-11-2018 15:00-00 | -- :- | 2 | Sensors disconnected during rewiring of post box | TSoil20, TSoil50, |
04-12-2018 00:00-05 | 04-12-2018 12:15-00 | 2 | Sensor disconnected for testing | TSoil30, TSoil100, |
26-11-2018 00:05-00 | 26-11-2018 14:48-00 | 2 | Sensor disconnected for testing | TSoil10, |
26-11-2018 09:30-00 | 26-11-2018 11:00-00 | 2 | Sensors disconnected during rewiring of post box | Tgrass, Tconc, Tsoil, G, |
21-11-2018 16:00-00 | 04-12-2018 12:15-00 | 2 | Suspect readings from all soil temperature probes | TSoil5, TSoil20, TSoil30, TSoil50, TSoil100, |
21-11-2018 15:00-00 | 21-11-2018 16:00-00 | 2 | Sensors disconnected for testing | TSoil5, TSoil20, TSoil30, TSoil50, TSoil100, |
19-11-2018 00:00-00 | 26-11-2018 00:00-00 | 2 | Sensor disconnected for testing | TSoil10, |
02-10-2018 09:00-00 | 20-10-2018 00:00-00 | 2 | Soil temperature sensor disconected during ground works | Tsoil, |
25-07-2018 08:40-00 | 25-07-2018 08:55-00 | 2 | Disconnection of sensor for cleaning | PG, |
23-07-2018 00:00-00 | 23-07-2018 14:15-00 | 2 | Replacement of surface temperature probes and amps | Tgrass, Tconc, Tsoil, |
03-07-2018 09:03-00 | 03-07-2018 09:14-00 | 2 | Suntracker routine cleaning maintenance | Sd, Sb, |
06-06-2018 10:17-00 | 06-06-2018 10:33-00 | 2 | Suntracker routine cleaning maintenance | Sd, Sb, |
30-05-2018 13:30-00 | 30-05-2018 14:30-00 | 1 | Grass cutting | *, |
23-05-2018 13:26-00 | 23-05-2018 13:38-00 | 1 | 6m mast maintainance | T4.48, U6.4, U3.2, U4.48, |
09-05-2018 10:28-00 | 09-05-2018 10:53-00 | 2 | Sun tracker routine cleaning / maintenance | Sd, Sb, |
08-05-2018 14:14-00 | 08-05-2018 14:14-30 | 2 | Rain gauge affected by grass cutting and blowing | Rain, |
08-05-2018 13:20-00 | 08-05-2018 15:20-00 | 1 | Grass cutting | *, |
02-05-2018 00:05-00 | 02-05-2018 11:30-00 | 2 | Pulse channel cards not available on logger | U2, U5, U10, U0.56, U0.8, U1.12, U1.6, U2.24, U3.2, U4.48, U6.4, Rain, |
30-04-2018 16:20-00 | 02-05-2018 11:30-00 | 2 | Pulse channel cards not available on logger | U2, U5, U10, U0.56, U0.8, U1.12, U1.6, U2.24, U3.2, U4.48, U6.4, Rain, |
30-04-2018 14:40-00 | 30-04-2018 16:20-00 | 2 | Data logger swapped for loan unit from Campbell | *, |
19-03-2018 00:00-00 | 27-03-2018 10:40-00 | 2 | Tsoil PRT disconnected following failure in snow | Tsoil, |
22-11-2017 10:26-25 | 22-11-2017 10:53-05 | 2 | Druck DPI140 pressure transducer disconnected for electrical safetry testing | P, |
15-11-2017 00:00-00 | -- :- | 2 | Flat plate level 1 data disabled | iFP, |
31-10-2017 14:00-00 | 31-10-2017 15:00-00 | 1 | Grass cutting | *, |
25-10-2017 08:00-00 | 27-10-2017 10:00-00 | 2 | Removal of soil temperature probe electronics and cabling for refurbishment | TSoil5, TSoil10, TSoil20, TSoil30, TSoil50, TSoil100, |
23-10-2017 00:00-00 | -- :- | 1 | All data channels at risk due to testing and data logger software changes | *, |
11-10-2017 13:15-00 | 11-10-2017 14:45-00 | 1 | Grass cutting | *, |
03-10-2017 14:50-00 | 03-10-2017 15:15-00 | 2 | flat plate GDACCs moved for strimming | iFP, |
03-10-2017 14:00-00 | 03-10-2017 15:25-00 | 1 | Grass cutting | *, |
15-08-2017 08:30-00 | 15-08-2017 14:00-00 | 1 | Grass cutting | *, |
14-08-2017 12:11-00 | 14-08-2017 12:15-00 | 2 | Field mill replaced with spare unit | PG, |
14-08-2017 08:44-00 | 14-08-2017 12:11-00 | 2 | Disconnection of sensor for checking | PG, |
28-06-2017 11:42-30 | 28-06-2017 11:53-00 | 2 | Disconnection of sensors for checking | TSoil5, TSoil10, |
21-06-2017 09:30-00 | -- :- | 1 | Possible problem with sensor | TSoil10, |
20-06-2017 21:02-10 | 20-06-2017 23:47-59 | 2 | Unknown sensor problem | P, |
16-06-2017 09:05-00 | 16-06-2017 10:16-00 | 2 | Ceilometer disconnected to test connections on Lightning PC for UAEREP. 19200 baud. | |
12-06-2017 08:00-00 | 12-06-2017 11:00-00 | 1 | Grass cutting by grounds department | *, |
22-05-2017 00:00-01 | 22-05-2017 09:18-00 | 2 | James new corona sensor installed for testing | JamesCoronaTest, |
27-04-2017 10:30-46 | 27-04-2017 13:25-25 | 1 | Sonic and licor disconnected while sonic changed for newly calibrated serial number 101703 | |
25-04-2017 12:35-00 | 25-04-2017 13:44-00 | 1 | Grass cutting | *, |
19-04-2017 09:30-00 | 19-04-2017 12:30-00 | 1 | Grass cutting | *, |
28-03-2017 12:30-00 | 28-03-2017 12:55-00 | 1 | Grass cutting | *, |
27-03-2017 12:30-00 | 27-03-2017 14:00-00 | 1 | Grass cutting | *, |
27-03-2017 13:15-00 | 27-03-2017 13:25-00 | 2 | cleaning of flat plate GDACCs | iFP, |
16-03-2017 09:09-00 | 16-03-2017 09:11-00 | 0 | Wick changed | Tw, |
09-03-2017 14:50-00 | 09-03-2017 15:35-00 | 1 | Gill mast lowered for maintenance | U, W, |
09-03-2017 14:00-00 | 10-03-2017 11:25-00 | 1 | Gill anemometer very noisy and keeps stopping. | V, |
01-03-2017 09:17-00 | 01-03-2017 09:19-40 | 2 | Tsoil 10 cabling and amp checked for water ingress | TSoil10, |
20-02-2017 09:35-42 | 20-02-2017 14:30-00 | 1 | Mast tipped for temp profile testing | T1.12, T0.56, T2.24, T4.48, U6.4, U3.2, U4.48, |
13-02-2017 13:40-00 | 13-02-2017 13:50-00 | 2 | Field mill swapped for original unit | PG, |
14-02-2017 13:45-00 | 14-02-2017 14:40-00 | 1 | Ice / water zero check on T1.12 | T1.12, |
13-02-2017 13:50-00 | 13-02-2017 14:30-00 | 2 | Field mill removed for checking and replaced with spare unit | PG, |
13-02-2017 11:45-00 | 13-02-2017 12:00-00 | 1 | Cable check on T0.56 | T0.56, |
13-02-2017 10:20-00 | 13-02-2017 10:30-00 | 1 | Ice / water zero check on T1.12 | T1.12, |
13-02-2017 10:35-00 | 13-02-2017 10:45-00 | 1 | Ice / water zero check on Td | Td, |
01-12-2016 13:50-00 | -- :- | 2 | GDACCs disconected for maintenance | iCP, |
01-12-2016 10:00-00 | 08-02-2017 13:00-00 | 2 | 10m tower failure | U5, U10, Dirn10, Sdur, Sb(csd3), |
06-07-2016 10:00-00 | 06-07-2016 11:15-00 | 1 | 10m tower tilted for maintenance | U5, U10, Dirn10, |
06-07-2016 10:00-00 | 06-07-2016 11:15-00 | 2 | CSD3 relocated to 10m tower | SdurTest, |
27-06-2016 00:00-00 | 27-06-2016 00:00-00 | 2 | CSD1 (Met Office) sunshine duration recorder removed for return | SdurMOTest, SbMOTest, |
22-06-2016 09:15-00 | 22-06-2016 09:25-00 | 1 | New wick installed on Twet | Tw, |
23-06-2016 00:00-00 | 24-06-2016 00:00-00 | 2 | Major sun tracker failure | Sb, Sd, |
27-05-2016 13:40-00 | 27-05-2016 15:40-00 | 2 | Power outage | *, |
26-05-2016 00:00-00 | 02-06-2016 10:30-00 | 2 | Instrument maintenance | iCP, |
09-05-2016 00:00-00 | 12-05-2016 00:00-00 | 2 | Suntracker maintenance | Sg, |
05-05-2016 00:00-00 | 23-05-2016 00:00-00 | 2 | Suntracker maintenance | Sd, Sb, |
27-04-2016 08:50-00 | 27-04-2016 09:05-00 | 2 | Disconnection of field mill to lower instrument to 1.5m | PGAngleTest, |
18-04-2016 00:00-00 | 20-04-2016 12:00-00 | 1 | Suntacker out of alignment | Sd, Sb, |
19-04-2016 14:15-00 | 19-04-2016 14:45-00 | 2 | Connector maintenance on CSD3 | SdurTest, Sb(csd3), |
17-04-2016 08:40-00 | 17-04-2016 10:50-00 | 2 | Power outage | *, |
31-03-2016 13:40-00 | 31-03-2016 13:50-00 | 2 | 2m anemometer replaced | U2, |
31-03-2016 13:40-00 | 31-03-2016 14:20-00 | 1 | 10m mast lowered for anemometer change | U5, U10, Dirn10, |
15-03-2016 00:00-00 | 15-03-2016 00:00-00 | 0 | Calibration of NRLite changed (calibrated against CNR4) | Rn, |
07-03-2016 14:00-00 | 07-03-2016 14:15-00 | 1 | Disruptions to mains power in hut | *, |
07-03-2016 10:50-00 | 07-03-2016 16:10-00 | 1 | Sun tracker returned to home position and switched off for mains supply maintenance | Sd, Sb, |
03-03-2016 15:30-00 | 03-03-2016 15:35-00 | 1 | Changes to power supply cables | *, |
03-03-2016 00:00-00 | 03-03-2016 00:00-00 | 2 | Vaisala pressure transducer removed from Ch1 | ch1, |
29-02-2016 09:00-00 | 29-02-2016 11:45-00 | 0 | Installation of field mill on post 12 for RGH research | ch55, PGAngleTest, |
11-02-2016 09:15-00 | 11-02-2016 10:55-00 | 1 | 10m mast lowered & raised for temporary student project installation | Dirn12, U5, U10, |
28-01-2016 15:25-00 | 28-01-2016 15:35-00 | 1 | 10m mast lowered & raised for temporary student project installation | Dirn11, U5, U10, |
28-01-2016 09:30-00 | 28-01-2016 09:45-00 | 1 | 10m mast lowered & raised for temporary student project installation | Dirn10, U5, U10, |
05-11-2015 13:55-00 | 03-30-2016 00:00-00 | 0 | Druck DPI140 added for testing pending permanent connection | PDruckTest, |
02-11-2015 15:20-00 | 19-01-2016 00:00-00 | 0 | Fog sensor (RGH test) added | FogSensorOrangeTest, FogSensorBlueTest, FogSensorPurpleTest, |
19-10-2015 10:40-00 | 28-10-2015 10:05-00 | 2 | equivilant +30 deg C resistor connected to PRT amp for test | Tsoil50Test, |
12-10-2015 13:50-00 | 19-10-2015 10:35-00 | 1 | 2.5V reference replaced sensor for test | Tsoil50Test, |
06-10-2015 14:00-00 | 12-10-2015 13:40-00 | 0 | 2.5V reference added on post 5 cable 5 | |
09-09-2015 14:50-00 | 09-09-2015 15:20-00 | 1 | 10m mast lowered for maintenance | U5, U10, Dirn10, |
30-08-2015 08:45-00 | 16-03-2016 10:30-00 | 2 | Instrument removed for calibration | Sdur, Sb(csd3), |
28-08-2015 16:00-00 | 01-09-2015 11:00-00 | 1 | Electrical works on Campus - data may be suspect or interrupted | *, |
26-08-2015 10:00-00 | 26-08-2015 10:45-00 | 1 | UPS brought into service | *, |
21-08-2015 08:30-00 | 21-08-2015 09:00-00 | 1 | Profile mast lowered for maintenance | U0.56, U0.8, U1.12, U1.6, U2.24, U3.2, U4.48, U6.4, |
20-08-2015 12:45-00 | 20-08-2015 13:30-00 | 1 | Profile mast lowered for maintenance | U0.56, U0.8, U1.12, U1.6, U2.24, U3.2, U4.48, U6.4, |
17-08-2015 10:45-00 | 14-09-2015 00:00-00 | 2 | Instrument removed for maintenance | PointDischarge, |
17-08-2015 09:00-00 | 01-09-2015 12:00-00 | 2 | New instruments added - still under test | T0.56Test, T1.12Test, T2.24Test, T4.48Test, |
13-08-2015 09:10-00 | 13-08-2015 09:30-00 | 0 | New instruments added | SdurMOTest, SbMOTest, |
13-08-2015 09:10-00 | 13-08-2015 09:30-00 | 1 | Instrument powered down temporarily | Sdur, Sb(csd3), |
07-08-2015 14:45-00 | 07-08-2015 14:50-00 | 1 | Instrument powered down temporarily | P, |
07-08-2015 00:00-00 | 07-08-2015 00:10-00 | 0 | Wind profile added | U0.56, U0.8, U1.12, U1.6, U2.24, U3.2, U4.48, U6.4, |
06-08-2015 12:15-00 | 06-08-2015 13:15-00 | 1 | Grass cutting | *, |
02-08-2015 07:30-00 | 02-08-2015 07:50-00 | 1 | New wick installed | Tw, |
27-07-2015 17:00-00 | 28-07-2015 09:30-00 | 1 | Planned system maintenance | Sb(csd3), |
27-07-2015 00:00-00 | 27-07-2015 17:00-00 | 1 | Planned system maintenance | *, |
23-07-2015 09:00-00 | 23-07-2015 15:00-00 | 0 | Building new 6m mast | *, |
22-07-2015 09:00-00 | 22-07-2015 16:00-00 | 0 | Building new 6m mast | *, |
15-07-2015 11:30-00 | 15-07-2015 12:30-00 | 1 | Sensor disturbed | U2, |
07-07-2015 11:55-00 | 07-07-2015 12:00-00 | 1 | Sensor disconnected | U2, |
02-07-2015 00:00-00 | 02-07-2015 13:10-00 | 1 | Sensor recalibration | Rn, |
29-06-2015 11:00-00 | 29-06-2015 16:00-00 | 1 | Grass cutting | *, |
29-06-2015 08:00-00 | 01-07-2015 00:00-00 | 1 | Started building new 6m mast | PG, iCP, iFP, R38, |
28-06-2015 00:00-00 | 28-06-2015 09:00-00 | 1 | Instrument disturbed by wildlife | iFP, |
20-06-2015 00:00-00 | 21-06-2015 09:07-00 | 1 | Instrument disturbed by wildlife | iFP, |
12-06-2015 08:00-00 | 12-06-2015 09:36-00 | 2 | LoggerNet comms lost. Logger rebooted | *, |
27-05-2015 10:20-00 | 28-05-2015 00:00-00 | 1 | Sensor recalibration | Sg, Sdw, |
14-05-2015 13:00-00 | 14-05-2015 17:00-00 | 0 | New instruments added | SuwTest, SdwTest, LuwTest, LdwTest, |
14-05-2015 00:00-00 | 14-05-2015 08:20-00 | 1 | Radiometer amp calibration changed at 08:20. Previous data erroneous | Sg, |
30-04-2015 09:10-00 | 30-04-2015 09:30-00 | 2 | Logger shutdown for upgrade | *, |
27-04-2015 00:00-00 | 28-04-2015 00:00-00 | 1 | Suntracker problems, instruments disturbed | Sd, Sg, |
16-04-2015 00:00-00 | 23-04-2015 00:00-00 | 1 | Sensor recalibration | Dirn10, |
15-04-2015 08:00-00 | 15-04-2015 10:00-00 | 1 | Suntracker problems, instruments disturbed | Sg, Sdw, |
15-04-2015 08:30-00 | 16-04-2015 08:30-00 | 1 | Maintenance and repairs | U2, U5, U10, Dirn10, |
14-04-2015 15:10-00 | 14-04-2015 15:40-00 | 1 | Suntracker problems, instruments disturbed | Sd, Sg, |
14-04-2015 00:00-00 | 14-04-2015 15:10-00 | 1 | Suntacker re-aligned | Sd, Sg, |
14-04-2015 00:00-00 | 01-05-2015 00:00-00 | 1 | Suntracker problems, instrument disabled | Sb, |
13-04-2015 00:00-00 | 13-04-2015 14:00-00 | 1 | Instrument maintenance | Tsoil30, Tsoil50, Tsoil100, |
13-03-2015 08:00-00 | 22-03-2015 00:00-00 | 0 | Teclipse added to ch 26 | Teclipse, |
05-01-2015 00:00-00 | 06-01-2015 00:00-00 | 1 | Instrument maintenance | Tsoil10, |
17-02-2014 01:00-00 | 23-04-2015 00:00-00 | 1 | Spurious data spikes - unexplained | Dirn10, |
24-09-2014 00:00-00 | 14-05-2015 13:00-00 | 2 | Data not viable, instruments under test | Suw, Sdw, |
02-04-2012 17:20-00 | 02-04-2012 17:25-00 | 2 | System shutdown to configure PC for new RDG-HOME\sws09a logon | *, |
02-04-2012 16:35-00 | 02-04-2012 16:55-00 | 2 | System shutdown to configure PC for new RDG-HOME\sws09a logon | *, |
30-03-2012 11:45-00 | 30-03-2012 12:05-00 | 2 | System shutdown while UPS installed and tested | *, |
29-03-2012 14:00-00 | 29-03-2012 16:55-00 | 2 | System shutdown for relocation to workshop during new hut rebuild | *, |
28-03-2012 16:05-00 | 29-03-2012 12:00-00 | 1 | New diff channel converters noisy and with offsets. Old converters reinstalled as found to be compatible with new interconnect cables | *, |
28-03-2012 11:50-00 | 28-03-2012 16:05-00 | 2 | System shutdown during data copy to MET-DAS II PC | *, |
26-03-2012 09:55-00 | 26-03-2012 11:35-00 | 2 | Re-routing cables ready for hut rebuild caused temporary data loss | Evap, R11, |
26-03-2012 09:55-00 | 29-03-2012 12:10-00 | 2 | Re-routing cables ready for hut rebuild, Druck DPI140 disconnected | P, |
22-03-2012 14:15-00 | 22-03-2012 15:15-00 | 2 | Re-routing cables ready for hut rebuild caused temporary data loss | *, |
22-03-2012 14:00-00 | 23-03-2012 15:15-00 | 2 | Re-routing cables ready for hut rebuild caused temporary data loss | U10, Dirn10, R12, R13, R14, R15, |
22-03-2012 14:00-00 | 22-03-2012 16:00-00 | 1 | Re-routing cables ready for hut rebuild caused temporary data loss | Tdry, Twet, RH, Tdew, |
02-02-2012 10:00-00 | 02-02-2012 11:00-00 | 1 | Temporary loss of power to post 3 cable 1 resulted in a data spike | Tdry, Twet, |
01-12-2011 11:00-00 | 30-04-2012 14:40-00 | 1 | bi-log amp cabling moved from Post 10 to Post 11. Instrument stays co-located with Field Mill. Second output patched into ch24 for testing. | R11, Evap, |
14-10-2011 10:00-00 | 14-10-2011 15:30-00 | 1 | Signal cables re-routed as part of RUAO buildings upgrade | *, |
10-10-2011 09:20-00 | 10-10-2011 23:55-00 | 1 | Signal cables re-routed as part of RUAO buildings upgrade | *, |
03-10-2011 15:30-00 | 03-10-2011 16:00-00 | 1 | instruments removed and cleaned and heaters upgraded | iCP, iFP, |
28-09-2011 14:10-00 | 28-09-2011 14:30-00 | 1 | Gauze cleaned | Rain, |
22-09-2011 11:30-00 | 22-09-2011 12:15-00 | 2 | System shutdown for annual electrical safety testing | *, |
02-09-2011 14:45-00 | 02-10-2011 14:45-00 | 2 | Channel 14 disabled until signalling problems can be investigated | Tsoil10, |
31-08-2011 12:00-00 | 31-08-2011 12:30-00 | 1 | power disconnected during installation of bilog amp | PG, |
31-08-2011 12:00-00 | 31-08-2011 12:30-00 | 1 | bi-log amp installed for testing | R11, |
08-08-2011 18:55-00 | 17-08-2011 14:20-00 | 1 | new fwprt (number 123) installed in cage | Soil moist50, |
07-06-2011 09:00-00 | 15-08-2011 15:45-00 | 1 | Power supply (black - neutral) loose in post box. Re-connected and applied WD40 | Tsoil10, |
29-07-2011 14:10-00 | 15-08-2011 09:00-00 | 1 | Student project involves variable ventilation of screen | Tdry, Twet, RH, |
28-07-2011 09:15-00 | 28-07-2011 09:30-00 | 1 | Installing fans in Stevenson screen | Tdry, Twet, RH, |
01-07-2011 14:30-00 | 01-07-2011 23:55-00 | 1 | Cable shortened and run through new underground conduit | Rain, |
24-06-2011 14:00-00 | 24-06-2011 15:25-00 | 1 | Buccaneer connectors on cable to Tair very corroded. Replaced those, and part of cable to Sdtrk | Tair, Sdtrk, |
15-06-2011 13:50-00 | 15-06-2011 14:05-00 | 1 | Raingauge moved to new position 3.5m south of post 5 | Rain, |
27-05-2011 09:20-00 | 27-05-2011 10:15-00 | 1 | Amps moved during fault diagnostics on Tsoil50 | Tsoil5, Tsoil10, Tsoil20, Tsoil30, Tsoil100, |
26-05-2011 17:45-00 | 27-05-2011 10:15-00 | 1 | Still unknown.Replaced amp | Tsoil50, |
22-05-2011 23:15-00 | 25-05-2011 14:20-00 | 1 | Still unknown. Put WD40 on 6-way buccaneer going into post 6 | Tsoil50, |
19-05-2011 14:20-00 | 19-05-2011 14:55-00 | 1 | Sensors disturbed while adding spi-wrap to cables | Tconc, Tsoil, Tgrass, |
18-05-2011 11:10-00 | 18-05-2011 12:20-00 | 1 | Signals disturbed while investigating Tsoil50 | Tconc, Tsoil, Tgrass, Tsoil5, Tsoil10, Tsoil20, Tsoil30, Tsoil50, Tsoil100, G, |
10-05-2011 21:00-00 | 18-05-2011 12:20-00 | 1 | Still unknown. Connections? | Tsoil50, |
13-05-2011 11:30-00 | 13-05-2011 11:40-00 | 1 | Tconc moved during final cable routeing | Tconc, |
11-05-2011 09:45-00 | 11-05-2011 09:50-00 | 1 | Wick changed and reservoir cleaned | Twet, |
09-05-2011 11:35-00 | 09-05-2011 15:50-00 | 1 | Disconnected for pit maintenance | iCP, iFP, |
03-05-2011 15:40-00 | 15-05-2011 14:00-00 | 1 | AspWet wick removed for testing | AspWet, |
27-04-2011 12:00-00 | 03-05-2011 15:40-00 | 1 | Installation of Delta T aspirated psychrometer for student project. AspWet is currently not wetted. | AspDry, AspWet, |
19-04-2011 11:40-00 | 03-05-2011 15:20-00 | 1 | probably suspect PRT amplifier. Amp replaced | Tconc, |
13-04-2011 10:10-00 | 13-04-2011 11:10-00 | 1 | All connections remade | Tsoil50, |
11-04-2011 23:00-00 | 13-04-2011 10:10-00 | 1 | Still unknown. Connections? | Tsoil50, |
11-04-2011 14:35-00 | 11-04-2011 16:00-00 | 2 | Instrument cables shortened for final installation | Tsoil5, Tsoil10, Tsoil20, Tsoil30, Tsoil50, Tsoil100, |
08-04-2011 23:00-00 | 11-04-2011 10:30-00 | 1 | Still unknown. Connections? | Tsoil50, |
08-04-2011 09:35-00 | 08-04-2011 11:35-00 | 1 | Vaisala soil temperature probes installed in new pit | Tsoil5, Tsoil10, Tsoil20, Tsoil30, Tsoil50, Tsoil100, |
21-03-2011 12:00-00 | -- :- | 0 | added sonde radiometer thermistor (MORSE) for testing in fine weather | R5, |
21-03-2011 12:00-00 | 21-03-2011 12:05-00 | 0 | removed gieger input | R5, |
17-03-2011 15:10-00 | 17-03-2011 15:30-00 | 1 | cables disconnected during installation of new instrument and associated rewiring of post box | Rain, Lin, Lout, |
17-03-2011 15:40-00 | -- :- | 0 | added sonde radiometer (MORSE) for testing in fine weather | R6, |
17-12-2010 10:15-00 | 17-12-2010 10:35-00 | 2 | system shutdown and restart | *, |
01-11-2010 12:10-00 | 13-12-2010 14:20-00 | 2 | Input used for second Lin from adjacent instrument for comparison | Lout, |
19-10-2010 13:55-00 | 21-10-2010 15:30-00 | 2 | Instrument removed for output testing | U2, Urun2, |
15-10-2010 13:55-00 | 15-10-2010 15:50-00 | 2 | Instrument removed and replaced with pulse outuput type for comparison with 'obs' 2m anemometer | U2, Urun2, |
11-10-2010 10:15-00 | 15-10-2010 15:50-00 | 1 | U2 height raised during testing of manual wind run counter | U2, Urun2, |
03-09-2010 10:10-00 | 03-09-2010 12:30-00 | 2 | Disk management | *, |
28-08-2010 19:20-00 | 02-09-2010 10:00-00 | 2 | Unexplained system shutdown. Another power outage? | *, |
28-08-2010 16:00-00 | 28-08-2010 17:45-00 | 2 | Mains power to Earley gate swithed off by SEC for maintenance | *, |
26-08-2010 10:00-00 | 26-08-2010 10:25-00 | 1 | GDACSS replaced | R9, R10, |
17-08-2010 11:00-00 | 18-08-2010 11:30-00 | 2 | system stopped logging with possible NI DMA error message | *, |
28-07-2010 15:15-00 | 26-08-2010 10:00-00 | 2 | testing of ISRAEL GDACCS | Tsoil5, Tsoil10, Tsoil20, Tsoil30, Tsoil50, Tsoil100, |
27-07-2010 11:30-00 | 27-07-2010 16:45-00 | 2 | maintenance of mains supply to tower instruments | Sonic Dirn, Sonic Mag, Sonic W, |
27-07-2010 08:40-00 | 27-07-2010 14:45-00 | 2 | temporary test of JCI131 | G, |
20-07-2010 08:35-00 | 20-07-2010 08:45-00 | 1 | Wet bulb wick changed | Twet, |
19-07-2010 08:35-00 | 19-07-2010 08:45-00 | 1 | Wet bulb wick wetted | Twet, |
11-07-2010 08:25-00 | 12-07-2010 10:45-00 | 2 | System shut down and restarted and seems OK | *, |
28-06-2010 12:00-00 | 28-06-2010 12:20-00 | 1 | Ch2 (Sd) now cabled from suntracker (old Sdtrk) rather than post 5. Old Sd now on ch33 called Sd_band | Sd, Sdtrk, Sd_band, |
22-06-2010 14:50-00 | 22-06-2010 15:15-00 | 1 | Investigation of wiring pin-outs on diff amp connector. | R1, R2, R3, R4, R5, R6, R7, R8, R9, R10, R11, R12, R13, R14, R15, R16, |
22-06-2010 09:30-00 | 22-06-2010 10:15-00 | 1 | Power to PG interrupted during installation of research instrument | PG, |
31-05-2010 00:00-00 | 31-05-2010 00:05-00 | 2 | Geiger 1 removed following completion of MSc project | R6, |
14-05-2010 15:20-00 | 14-05-2010 16:00-00 | 1 | Analogue output from geigers 3 & 1 now connected to R5 & R6 respectively. Patch cabling changed to post cable 13 from cable 11 | R3, R4, R5, R6, |
23-04-2010 10:15-00 | 28-06-2010 12:20-00 | 2 | Sd currently disabled. To be replaced shortly with Sdtrk (requires channel swapping!) | Sd, |
16-04-2010 14:15-00 | 16-04-2010 14:30-00 | 1 | Geiger 2 replaced in original position. Geiger 1 added to correlator with 2 and 3 | R3, R4, |
14-04-2010 12:25-00 | 16-04-2010 14:15-00 | 1 | Straight swap of Geiger 1 in place of Geiger 2 using all the same connections. Testing Geiger 1 | R3, R4, |
30-03-2010 14:20-00 | 30-03-2010 14:55-00 | 1 | Straight swap of Geiger 1 in place of Geiger 2 using all the same connections. Testing Geiger 1 | R3, R4, |
29-03-2010 14:10-00 | 29-03-2010 15:00-00 | 1 | Straight swap of Geiger 1 in place of Geiger 2 using all the same connections. Testing Geiger 1 | R3, R4, |
23-03-2010 14:00-00 | 24-03-2010 00:00-00 | 1 | New GDACSS calibration from 00.00 on day 83 | R9, R10, |
18-03-2010 12:40-00 | 18-03-2010 16:25-00 | 1 | Straight swap of Geiger 1 in place of Geiger 2 using all the same connections. Testing Geiger 1 | R3, R4, |
18-03-2010 12:00-00 | 23-03-2010 14:00-00 | 2 | GDACSS removed for repair and recalibration | R9, R10, |
10-02-2010 18:00-00 | 11-02-2010 12:10-00 | 1 | Wick had dried out. Re-wetted 42, 12.10 | Twet, |
08-02-2010 00:00-00 | 08-02-2010 16:15-00 | 1 | Change of screen temperature sensor calibration. Previously processed data incorrect from 00.00 to 16.15 | Tdry, Twet, |
08-02-2010 14:25-00 | 08-02-2010 16:15-00 | 1 | Installation of new sensors and test of calibration in ice/water etc | Tdry, Twet, |
01-02-2010 13:55-00 | 01-02-2010 16:25-00 | 1 | Tdry and Twet PRT sensors placed in ice/water for 0 deg check | Tdry, Twet, |
18-01-2010 14:00-00 | 21-01-2010 15:10-00 | 1 | Fault finding and repair of faulty Lin (CNR1) | Lout, Sout, |
18-01-2010 15:00-00 | 18-01-2010 15:35-00 | 1 | Tdry and Twet PRT sensors placed in ice/water for 0 deg check | Tdry, Twet, |
18-01-2010 14:35-00 | 21-01-2010 15:15-00 | 2 | Lin disabled while spurious readings investigated (repaired 1510 on day 21) | Lin, Tsky, |
22-12-2009 12:35-00 | 22-12-2009 12:55-00 | 1 | Installation of new tracker patch box and wiring | Sg, Sdtrk, Sd, |
15-12-2009 16:15-00 | 16-12-2009 10:20-00 | 1 | No supply connected to newly wired Druck pressure sensor amplifier | P, |
14-12-2009 16:50-00 | 15-12-2009 16:15-00 | 1 | Continuation of major re-wiring of internal patch boxes | *, |
14-12-2009 09:50-00 | 14-12-2009 16:50-00 | 2 | Major re-wiring of internal patch boxes | *, |
09-12-2009 11:05-00 | 09-12-2009 11:20-00 | 2 | Temporary removal of sensor power during replacement of 12V power supply | *, |
08-12-2009 12:30-00 | 08-12-2009 23:55-00 | 0 | Installed IAR-N on chR15 and IAR-NT on chR16 at standard height on post 10 | R15, R16, |
07-12-2009 13:55-00 | 07-12-2009 15:00-00 | 1 | Tested geiger 1. Geigers 2 & 3 and correlator moved to FRL | R3, R4, |
07-12-2009 13:55-00 | 07-12-2009 14:00-00 | 2 | Geiger 3 analogue out removed | R5, |
25-11-2009 09:45-00 | 25-11-2009 10:15-00 | 1 | installation of new (shorter) cable on 2m anemometer to post 4 | U2, |
24-11-2009 15:50-00 | 24-11-2009 16:20-00 | 2 | 2m anemometer and cabling moved to post 4. Anemometer moved approx 10m due E from previous location | U2, |
24-11-2009 10:20-00 | 24-11-2009 16:00-00 | 2 | Tdry Twet and RH sensor cabling moved to post 3. Sensor postions in screen remain unchanged | Tdry, Twet, RH, |
23-11-2009 15:20-00 | 23-11-2009 15:40-00 | 1 | Original PG now relocated to post 10 cable 2 (from post 4) | PG, |
17-11-2009 16:20-00 | 22-11-2009 15:40-00 | 0 | JCI 131 connected on post 10 for standardisation against PG on post 4 | R11, |
09-11-2009 16:00-00 | 17-11-2009 15:30-00 | 2 | disconnected for cable maintenance. New cabling installed underground between pit and new post 10. Heater supplies now from PSU via tower - post 10 link cables | R9, R10, |
16-09-2009 12:00-00 | 16-09-2009 12:15-00 | 1 | Wet bulb wick changed | Twet, |
16-09-2009 12:20-00 | 16-09-2009 15:40-00 | 2 | System shutdown for annual electrical safety testing | *, |
14-08-2009 14:50-00 | 14-08-2009 15:00-00 | 1 | Wet bulb wick changed | Twet, |
14-08-2009 09:30-00 | 14-08-2009 15:00-00 | 1 | Line driver on JCI40F removed | R6, |
13-08-2009 16:00-00 | 13-08-2009 16:10-00 | 0 | Wet bulb reservoir topped up | Twet, |
05-08-2009 16:00-00 | 05-08-2009 16:20-00 | 0 | JCI 140F installed at 1m height for calibration against PG | R6, |
04-08-2009 11:30-00 | 04-08-2009 11:40-00 | 1 | Geigers and correlator realigned on bench | R3, R4, R5, |
31-07-2009 09:30-00 | 31-07-2009 11:50-00 | 1 | Internal realignment of Geiger tubes. | R3, R4, R5, |
20-07-2009 12:00-00 | 20-07-2009 12:30-00 | 1 | Possible suspect / no data on all channels due to A/D cable connector problems | *, |
04-05-2009 00:00-00 | 05-05-2011 15:35-00 | 2 | Heat flux plate removed following failure.New plate to be installed early 2010 | G, |
17-07-2009 09:30-00 | 17-07-2009 09:40-00 | 1 | Geiger 2 relocated to top of Geiger 3 in NW corner of bench in FTL | R3, R4, R5, |
16-07-2009 14:00-00 | 16-07-2009 14:20-00 | 1 | Geiger 2 relocated (on longer lead) to NE corner of bench in FTL. Geiger 3 and correlator relocated to NW corner of bench in FTL | R3, R4, R5, |
06-07-2009 13:55-00 | 06-07-2009 14:00-00 | 1 | Wet bulb wick changed | Twet, |
19-06-2009 12:10-00 | 19-06-2009 12:30-00 | 2 | System reboot | *, |
19-06-2009 11:30-00 | 19-06-2009 11:40-00 | 2 | Ground temperature PRTs removed and disconnected | Tsoil5, Tsoil10, Tsoil20, Tsoil30, Tsoil50, Tsoil100, |
19-06-2009 11:30-00 | 19-06-2009 11:40-00 | 1 | Soil heat flux plate relocated 2m to E | G, |
01-06-2009 09:00-00 | 01-06-2009 11:40-00 | 1 | Wet bulb wick changed | Twet, |
14-05-2009 14:25-00 | 14-05-2009 14:30-00 | 0 | Wet bulb wick wetted | Twet, |
28-04-2009 09:50-00 | 29-04-2009 10:10-00 | 1 | New RH sensor head fitted | RH, |
18-02-2009 15:45-00 | 18-02-2009 16:15-00 | 1 | RH power supply interrupted during testing of field mill for DREAME project | RH, R6, |
18-02-2009 10:10-00 | 18-02-2009 11:00-00 | 1 | RH power supply interrupted during testing of field mill for DREAME project. DREAME field mill connected for testing | RH, R6, |
17-02-2009 12:20-00 | 17-02-2009 12:35-00 | 1 | RH power supply interrupted during testing of field mill for DREAME project. DREAME field mill connected for testing | RH, R6, |
12-02-2009 11:50-00 | 13-02-2009 12:35-00 | 2 | Tdry and Twet sensors removed for calibration | Tdry, Twet, |
11-02-2009 14:40-00 | 11-02-2009 15:15-00 | 2 | Sensor disconnected for calibration | Tdry, Twet, |
10-02-2009 11:40-00 | 10-02-2009 15:20-00 | 1 | no data due to calibration of geiger2 | R3, R4, R5, |
09-02-2009 09:50-00 | 10-02-2009 12:40-00 | 1 | Rainguage reset alarm clock reset to UTC from +1hr | Rain, |
15-01-2009 14:00-00 | 15-01-2009 14:20-00 | 1 | Twisted pair signal lead replaced with standard 6-core | R3, R4, R5, |
14-01-2009 14:25-00 | 14-01-2009 14:30-00 | 1 | moved from separate power supply to METFiDAS 14:25 | R3, R4, R5, |
01-01-2009 00:00-00 | 16-04-2009 14:30-00 | 1 | Sun tracking and attached instrument data suspect. Repaired 106. Valid data from 1320 | Sb, Sdtrk, |
10-12-2008 16:00-00 | 10-12-2008 16:20-00 | 0 | Geiger correlator installed on R3 Geiger 2 on R4 and Geiger 3 on R5 | R3, R4, R5, |
30-10-2008 09:15-00 | 30-10-2008 09:30-00 | 1 | Wet bulb wick replaced | Twet, |
27-10-2008 14:30-00 | 29-10-2008 15:00-00 | 1 | Wet bulb wick dried out. Manually wetted 303, 14.40 | Twet, |
29-09-2008 15:50-00 | 29-09-2008 16:20-00 | 1 | Corrugated and flat current plates moved at 16.00 into Pit | R9, R10, |
11-09-2008 12:40-00 | 11-09-2008 16:10-00 | 2 | System shutdown for electrical safety testing | *, |
01-09-2008 00:00-00 | 14-10-2008 13:40-00 | 1 | Wick had dried out. Re-wetted 11:26 | Twet, |
22-08-2008 10:35-00 | 17-09-2008 23:55-00 | 1 | Pgtest moved to post 4 for testing. Removed 261, 09.15 | R11, |
22-08-2008 10:35-00 | 22-08-2008 12:30-00 | 1 | Corrugated and flat current plates connected for testing. Pgtest moved to post 4 | R9, R10, R11, |
20-08-2008 11:20-00 | 20-08-2008 16:50-00 | 1 | Trials of different field mills taking place, data suspect from 11:20 to 16:50 | PG, R10, |
05-08-2008 14:20-00 | 19-08-2008 16:50-00 | 1 | Wick had dried out. Re-wetted 232, 10.00 | Twet, |
28-07-2008 16:20-00 | 20-08-2008 10:00-00 | 1 | field mill trials: Connected on post 4 (Ch42) for comparison against mill on post 1 (Ch29) | R10, |
28-07-2008 10:10-00 | 28-07-2008 11:20-00 | 2 | Logging PC rebooted | *, |
14-07-2008 12:20-00 | 14-07-2008 12:30-00 | 0 | RGH magnetometers disconnected | R10, R11, |
03-07-2008 11:15-00 | 03-07-2008 17:00-00 | 1 | Possible suspect data of all channels due to cable maintenance | *, |
01-07-2008 15:00-00 | 01-07-2008 15:25-00 | 1 | Possible suspect data of all channels due to cable maintenance | *, |
18-06-2008 10:10-00 | 18-06-2008 16:10-00 | 2 | Cable maintenance on 10m tower | U10, Dirn10, Sonic Dirn, Sonic Mag, Sonic W, |
17-06-2008 16:00-00 | 17-06-2008 16:10-00 | 2 | Cable maintenance on 10m tower | U10, Dirn10, Sonic Dirn, Sonic Mag, Sonic W, |
28-05-2008 13:00-00 | 10-06-2008 16:30-00 | 2 | Supply converter box removed for repair 13:00 - was full of water. | PG, |
28-05-2008 09:00-00 | 28-05-2008 10:20-00 | 1 | Data suspect | PG, |
20-05-2008 09:20-00 | 20-05-2008 13:00-00 | 1 | Shading ball on suntracker removed for maintenance | Sdtrk, |
17-05-2008 09:50-00 | 17-05-2008 12:45-00 | 1 | No DC supply to instrumentation during maintenance | *, |
07-05-2008 10:40-00 | 07-05-2008 10:50-00 | 0 | Tair installed 10.45 on R7 | R7, |
07-05-2008 10:25-00 | 07-05-2008 10:50-00 | 1 | Shade ring adjusted | Sd, |
29-04-2008 14:40-00 | 30-04-2008 10:35-00 | 2 | 10m tower istruments disconnected during work on cabling | U10, Dirn10, Sonic Dirn, Sonic Mag, Sonic W, |
19-04-2008 07:30-00 | 20-04-2008 11:00-00 | 2 | Mains supply failure - no data | *, |
03-04-2008 12:40-00 | 03-04-2008 15:10-00 | 1 | No DC supply to instrumentation during maintenance | Lin, PG, Lout, R1, R2, R3, R4, R5, R6, R7, R8, R9, R10, R11, R12, R13, R14, R15, R16, |
03-04-2008 08:15-00 | 03-04-2008 09:00-00 | 1 | No DC supply to instrumentation during maintenance | *, |
24-03-2008 09:09-00 | 31-03-2008 09:00-00 | 1 | Field mill supply converter box removed for repair (full of water!) | PG, |
20-01-2008 00:00-00 | 24-03-2008 12:00-00 | 1 | Data suspect | PG, |
-- :- | -- :- | |||
DD-MM-YYYY hh:mm-ss | DD-MM-YYYY hh:mm-ss | |||
Start date-- :- | End date-- :- | Code | Comments | Channels, |